OpenStudy (anonymous):

What are the possible number of negative zeros of f(x) = 2x7 – 2x6 + 7x5 – 7x4 – 4x3 + 4x2 ?

OpenStudy (anonymous):

hi u must factor it use this \[2x^7 – 2x^6 + 7x^5 – 7x^4 – 4x^3 + 4x^2=2x^6(x-1)+7x^4(x-1)-4x^2(x-1)\\=(x-1)(2x^6+7x^4-4x^2)\]

OpenStudy (anonymous):

Thank you! :D That was basically all I needed.

OpenStudy (anonymous):

welcome